OpenStudy (kirbykirby):

Integral:\[\int_0^{\infty}\frac{\cos t}{(t^2+1)^2}dt\] (Note: this might need methods of complex integration).

OpenStudy (anonymous):

I think first you have to convert cost somehow to a tan using sec^2+1=tan^2, then just use trig integral and do the bottom, something should be canceling out

OpenStudy (anonymous):

and Im sorry its sec^2-1=tan^2

OpenStudy (anonymous):

\[\int\limits_{0}^{n} \frac{ 1 }{ \sec(t) }*\frac{ 1 }{ (x^2+1)^2 }\] then you just do the \[x=1*\tan \theta\] \[dx = \sec^2(\theta) d \theta \] \[x^2+1=\sec^2(\theta)\] then you just substitute back everything

OpenStudy (anonymous):

@kirbykirby you know trig integral rule right?

OpenStudy (anonymous):

and that was 1/sec(x) not 1/sec(t) haha...

OpenStudy (kirbykirby):

But then it's something like \[ \int \frac{1}{\sec(\tan (\theta))}\frac{1}{\sec^2(\theta)}d\theta\]

OpenStudy (zarkon):

use residues

OpenStudy (anonymous):

I see... I didn't calculate it so I don't know that. But from something I find online http://memorize.com/half-angle-formulas it has something to do e^ix you can do some research on that